​Claim: The standard deviation of pulse rates of adult males is less than 12 bpm. For a random sample of 131 adult​ males, the pulse rates have a standard deviation of 11.8 bpm. Find the value of the test statistic.
Answer #1

Giren, n=131, 5=11.8, 60=12 Test statistics = (n-1) 3 = (131-1) (11.8) ? 122 =125.7027778